OpenStudy (anonymous):

can anyone provide any guidance on this series? Sum from n=0 to infinity of n((6^n)/n!)

OpenStudy (anonymous):

i calculated the first few terms and got 0+6+108+216+324+... and i recognize that if that first n was not there then it would be e^6 but im not sure what to do with the n

OpenStudy (anonymous):

haha hey again

OpenStudy (amistre64):

howdy :)

OpenStudy (amistre64):

e^6 is definantly part of it; do you know the answer already?

OpenStudy (anonymous):

lol well i know worlfram alpha says it should be 6e^6

OpenStudy (amistre64):

good, then the question is how to get there :)

OpenStudy (anonymous):

correct i thought maybe you can do the sums seperatly then mulitply together but that did not work out

OpenStudy (amistre64):

joes got a brilliant method im sure :)

OpenStudy (anonymous):

Do you know the infinite series for e^x?

OpenStudy (anonymous):

then i thought well i can end up getting n!/(n-1)! but that does not make sense for n=0

OpenStudy (anonymous):

yes i do its x^n/n!

OpenStudy (anonymous):

Take the infinite series for e^x:\[e^x=\sum_{n=0}^\infty \frac{x^n}{n!}\]Take the derivative of each side:\[\frac{d}{dx}(e^x)=\frac{d}{dx}\left(\sum_{n=0}^{\infty }\frac{x^n}{n!}\right)=\sum_{n=1}^{\infty }\frac{nx^{n-1}}{n!}\]

OpenStudy (anonymous):

This sorta matches what you want, but one thing is off.

OpenStudy (anonymous):

Right now we are at:\[e^x=\sum_{n=1}^{\infty}\frac{nx^{n-1}}{n!}\]The only thing that is off is the power. In your expression, you have a multiple of n, an n!, and an nth power. In mine, we have a multiple of n, an n!, and an (n-1)th power, so we need to fix that.

OpenStudy (anonymous):

by multiplying it by x right?

OpenStudy (anonymous):

thats correct :)

OpenStudy (anonymous):

so we end up with:\[xe^x=\sum_{n=1}^\infty \frac{nx^n}{n!}\]

OpenStudy (anonymous):

alright awesome so does it not matter that n goes from 1 to infinity as opposed to 0 to infinity?

OpenStudy (anonymous):

like how did you know to do the derivative and if you did that then to change the limits?

OpenStudy (amistre64):

at n=0 the add on is just zero right?

OpenStudy (anonymous):

if you plug in n=0, regardless of the value of x, we will end up with 0, so there is no difference between n=0 to infinity or n=1 to infinity.

OpenStudy (amistre64):

the derivative of e^x is equal to e^x

OpenStudy (anonymous):

I knew the change the limits in the summation because the first term was a constant (the constant 1), and the derivative would kill it.

OpenStudy (anonymous):

and the reason i choose to take a derivative in the first place is because your expression had an multiple of n in it, and derivatives bring down powers.

OpenStudy (anonymous):

alright cool thank you very much i get it now

OpenStudy (anonymous):

i really appreciate it